# Camera: Multi-Matrix Joint Compression for MoE Models via Micro-Expert Redundancy Analysis

### v3

@@ A.5 Mixed-precision Settings in Camera-Q
In the experiments in Table 2 of the main text, MC assigns different precisions across experts, whereas Camera-Q and Camera-Q apply varying precisions across micro-experts within each expert. Consequently, the weight matrix of each expert consists of multiple precision levels. We implement Camera-Q and its variants using a three-level mixed-precision scheme. To achieve an average bit-width of $b$, we define three precision levels: [b+1,b,b1][b+1,b,b-1] bits, with their respective proportions in the total weights set to [r,12r,r][r,1-2r,r]. For example, if the desired average precision is 2 bits with $r=0.1$, then the bit-widths $[3,2,1]$ are assigned in proportions $[0.1,0.8,0.1]$, yielding an average of $3\times 0.1+2\times 0.8+1\times 0.1=2.0$ bits. Since we perform group-wise quantization over every 128 vectors, each group stores an additional zero-point and scale factor, totaling $2\times 16=32$ bits (following the GPTQ algorithm (Frantar et al. 2022)). This adds an overhead of $32/128=0.25$ bit per weight, resulting in a total average cost of 2.25 bits per weight.
@@
A.6 Integrity of Micro-Experts
Joint compression of multiple matrices is a key innovation of our method, as it fully accounts for the functional correlation between weight matrices and the functional integrity of micro-experts. To illustrate this more clearly, we combine the Camera-P and Camera-Q working mechanism to discuss it.
Figure 9 illustrates the differences in weight processing across three algorithms, highlighting how expert weights are handled. All three methods share a common structure: the input hidden state $\mathbf{x}$ is first transformed into intermediate activations via $\mathbf{W}^{\mathrm{up}}$ and $\mathbf{W}^{\mathrm{gate}}$, followed by $\mathbf{W}^{\mathrm{down}}$ to produce the output $\mathbf{y}$. The red dashed box in the left subfigure highlights our cross-matrix definition of a micro-expert. Notably, each row of $\mathbf{W}^{\mathrm{up}}$ corresponds to a unique column in $\mathbf{W}^{\mathrm{down}}$, forming a one-to-one mapping of micro-experts. For better visualization in Figure 9, $\mathbf{W}^{\mathrm{up}}$ is drawn as its transpose, and all micro-experts are energy-ranked and rearranged so that more important ones appear to the left.
In the left subfigure, Camera-P prunes the less important tail micro-experts while preserving the functional integrity of the retained ones. In the center, Camera-Q allocates more bits to higher-energy micro-experts, ensuring that all weights within a given micro-expert share the same precision. In contrast, the right subfigure shows Camera-Q, which slices $\mathbf{W}^{\mathrm{up}}$ (and $\mathbf{W}^{\mathrm{gate}}$) along a direction orthogonal to that of Camera-Q. Each weight $\mathbf{w}^{\mathrm{up/gate}}$ in one micro-expert is quantized using all three precisions. In fact, most existing mixed-precision quantization methods, like Camera-Q, operate on individual matrices and assign precision along the input dimension (e.g., using the Hessian $\mathbf{H}=\mathbf{XX}^{\mathrm{T}}$ as an importance proxy). However, our analysis suggests that this single-matrix view neglects micro-expert integrity. The $\mathbf{w}^{\mathrm{up/gate}}$ components of high-energy micro-experts may receive inconsistent and insufficient precision guarantees. Our comparative experiments with Camera-Q further demonstrate that maintaining consistent precision within individual micro-experts leads to superior performance.

### v4

@@ A.5 Mixed-precision Settings in Camera-Q
In the experiments in Table 2 of the main text, MC assigns different precisions across experts, whereas Camera-Q and Camera-Q apply varying precisions across micro-experts within each expert. Consequently, the weight matrix of each expert consists of multiple precision levels. We implement Camera-Q and its variants using a three-level mixed-precision scheme. To achieve an average bit-width of $b$, we define three precision levels: [b+1,b,b1][b+1,b,b-1] bits, with their respective proportions in the total weights set to [r,12r,r][r,1-2r,r]. For example, if the desired average precision is 2 bits with $r=0.1$, then the bit-widths $[3,2,1]$ are assigned in proportions $[0.1,0.8,0.1]$, yielding an average of $3\times 0.1+2\times 0.8+1\times 0.1=2.0$ bits. Since we perform group-wise quantization over every 128 vectors, each group stores an additional zero-point and scale factor, totaling $2\times 16=32$ bits (following the GPTQ algorithm (gptq2022)). This adds an overhead of $32/128=0.25$ bit per weight, resulting in a total average cost of 2.25 bits per weight.